# FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models

### v3

@@ 3.2 Super-Resolution with Factorized Features
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
where $c^{\text{LR}}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
Our model comprises three main components: Coefficient Backbone $F_{\textbf{coeff}}$, Basis Swin Transformer $F_{\textbf{basis}}$, and Factorized Features Reconstruction. As shown in  Fig. 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations. Then, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq. 8, where $s$ is the scale factor.
We optimize model parameters using $L_{1}$ loss.
To demonstrate the effectiveness of our method, we use existing SR methods [16, 17, 91] as the Coefficient Backbone. For the Basis Swin Transformer, we employ simple Swin Transformer Blocks [59] with a series of special downsampling operations. Specifically, we use a dilation-like downsampling technique( Fig. 2.c) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
@@ 3.3 Image Compression with Factorized Features
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in  Sec. 1 and  Sec. 2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [39, 44], or decompression modules [75, 24], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through Factorized Features formulation as in  Eq. 7
Priors from SR. In addition, with our trained SR model described in  Sec. 3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
The overall pipeline is shown in  Fig. 6b. To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [55], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined as
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (9)
@@ 3.3.1 Multi-Image Compression
For each basis $b_{i}\in R^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ associated with any arbitrary image, we can consider it as encapsulating the inherent pixel structure. These bases can be combined into a unified generic basis that captures the structural distribution of images and potentially reduces noise. Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (10)
[…]
Since compression induces error, according to  Fig. 6b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error. To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients. as illustrated in Fig. 6c. This way, we can enjoy less information loss with basis while maintaining low bit rates.
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining [16]. Please refer to quantitative results in  Sec. 4.1 for details.
@@
SwinIR [48](ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD [92](CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT [17])(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT [18](ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT [15](CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L [15](CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
DAT-L* 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
@@
HAT-F-Basis-First (Ours) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
HAT-F-Concat (Ours) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
@@ 4.1 Image Super-Resolution
Experimental Setup. We conduct extensive experiments to validate the effectiveness of our Factorized Features representation for Super-Resolution tasks. Following the strategy outlined in [15] and [20], we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from [37], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [22], which is used for [15] and [20] pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR [48], HAT [16], DAT [17], and ATD [92]) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K [1] + Flickr2K [52]) as the finetuning dataset following [21, 15] for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 2e-4, batch size 32, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Features, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information. More implementation details can be found in supplementary materials.
Quantitative Results. Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [6], Set14 [90], BSD100 [63], Urban100 [33], and Manga109 [64]. For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [91], DAT [17], and HAT [16]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts, ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only [16] provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [14, 16]. The results demonstrate that its performance remains consistent with only minor perturbations.
@@
Furthermore, in traditional Fourier Series and other image processing methods [82], the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in  Fig. 3. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
@@
Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 7. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
@@
Visual Comparison. We provide the visual comparison in Fig. 7. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Multi-frame SR. Due to space limitations, we provide additional comparisons on multi-frame SR in the supplementary materials.
@@
Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
Tot Enc $\downarrow$ Tot Dec $\downarrow$
@@
Xie (MM’21) -0.78 2.93 6.00 50.0
Cheng (CVPR’22) 5.44 1.98 4.69 29.6
STF (CVPR’22) -4.31 0.14 0.13 99.9
ELIC (CVPR’22) -7.24 0.07 0.09 36.9
TCM (CVPR’23) -11.74 0.16 0.15 76.7
CCA (NeurIps’24) -20.97 0.11 0.09 64.9
@@
TCM-HAT-F-multi M=1 (Ours) 27.96 0.232 0.174 131.35
TCM-HAT-F-multi M=2 (Ours) 2.70 0.232 0.174 131.35
TCM-HAT-F-multi M=4 (Ours) -10.11 0.232 0.174 131.35
TCM-HAT-F-multi M=8 (Ours) -16.61 0.232 0.174 131.35
TCM-HAT-F-multi M=16 (Ours) -19.88 0.232 0.174 131.35
TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
Method Kodak CLIC Tecnick
[…]

### v4

@@ 4.2 Super-Resolution with Factorized Features
$\small\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl({\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl\{c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl\}\Bigl),$ (8)
where $c^{\text{LR}}(x)=F_{\textbf{coeff}}(I_{\text{LR}})(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$. Note that we sample the outputs $F_{\textbf{coeff}}(I_{\text{LR}})$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively. We omit convolution layers here for simiplicity.
Our model comprises three main components: Coefficient Backbone $F_{\textbf{coeff}}$, Basis Swin Transformer $F_{\textbf{basis}}$, and Factorized Features Reconstruction. As shown in  fig.˜5a, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations. Also, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  eq.˜8, where $s$ is the scale factor.
We optimize model parameters using $L_{1}$ loss. To be more clear, when the Basis Swin Transformer is optimized, the basis is optimized simultaneously, i.e. the models are trained to adapt to a content-aware basis, not calculated by heuristic rules. This design enables content-aware basis generation: smooth regions get low-frequency bases while textured areas receive high-frequency ones.
To demonstrate the effectiveness of our method, we use existing SR methods [16, 18, 109] as the Coefficient Backbone. For the Basis Swin Transformer, we employ simple Swin Transformer Blocks [70] with a series of special downsampling operations. Specifically, we use a dilation-like downsampling technique to accommodate the $\gamma$ sawtooth sampling pattern, where the details and visualization can be found in Supplementary Materials. The final basis is refined using additional upsampling and convolution layers.
@@ 4.3 Image Compression with Factorized Features
Image compression balances bit-rate and visual fidelity. Recent work recovers hidden details mainly via architectural tweaks—analysis transforms and entropy models [46, 51] or sophisticated decoders [88, 25]. We instead target the representation itself: our Factorized Features formulation (eq.˜7) explicitly models structural correlations, boosting quality.
Figure 5: Super-Resolution and Image Compression with Factorized Features. This figure illustrates how Factorized Features are used in Super-Resolution and Image Compression. (a) Given a low-resolution input image, we first extract $X_{coeff}$ feature with Coefficient Backbone. Next, we generate the basis and coefficient with Basis Swin Transformer and convolution layers, respectively, from the same $X_{coeff}$. Finally, the prediction is reconstructed by Factorized Features Reconstruction. (b) To decrease distortion in image compression, we replace the synthesis transform of the traditional learned image compression pipeline with (a) by aligning spatial resolution and latent channels.
Priors from SR. In addition, with our trained SR model described in  section˜4.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
The overall pipeline is shown in  fig.˜5b. To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [62], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in the Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with common loss function used in learned image compression, defined as
[…]
where $\hat{z}$ is the hyperprior, $R(\cdot)$ refers to the bit-rate cost, $D(\cdot)$ is the distortion term, and $\lambda$ controls the trade-off.
@@ 4.3.1 Multi-Image Compression
Intuitively, different images often share common frequency components. By learning a single, shared set of basis functions, we can leverage the frequently appearing similar or repetitive patterns across images to greatly reduce coding redundancy in multi-image compression. In practice, this means fewer total parameters and lower computational cost required when compressing an image collection.
Each per-image basis $b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ captures local pixel structure. We fuse the $M$ bases into a denoised, generic basis with a transformer $F_{\text{merge}}$ [83] applied at every position:
$\small b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (10)
Treating bases as tokens (prepended with a CLS token), the transformer’s output forms the merged basis.
Because compression noise corrupts the coefficient map $X_{\text{coeff}}$ (fig.˜5b), feeding it directly into the Basis Swin Transformer would amplify errors. Instead, we transmit the merged basis separately, alongside the quantized variables $^{Q(y,\mu),\hat{z}}$, and reconstruct each image with this clean basis plus its own decoded coefficients, yielding higher quality at similar bit-rates. A single- vs. multi-image compression comparison is in the supplementary materials.
Table 1: Quantitative comparisons on $4\times$ super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use the same-task pretraining [16], i.e., pretrained on ImageNet. Please refer to quantitative results in  section˜5.1 for details.
@@
EDSR [58](CVPR’17) 43.09 207.07 1,182 32.46 0.8968 28.80 0.7876 27.71 0.7420 26.64 0.8033 31.02 0.9148
RCAN  [111](ECCV’18) 15.59 65.52 1,176 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
SwinIR [55](ICCV’21) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
CAT-A+ [17](NIPS’22 Spotlight) 16.60 70.29 3,508 33.14 0.9059 29.23 0.7968 28.01 0.7516 27.99 0.8356 32.52 0.9293
ART  [108](ICLR’23 Spotlight) 16.56 69.94 3,010 33.04 0.9051 29.16 0.7958 27.97 0.7510 27.77 0.8321 32.31 0.9283
ATD [110](CVPR’24) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT [18])(ICCV’23) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT [19](ICLR’24) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
PFT [71](CVPR’25) 19.66 80.21 6,414 33.15 0.9065 29.29 0.7978 28.02 0.7527 28.20 0.8412 32.63 0.9306
HAT [15](CVPR’23) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
ATD-L 49.42 184.83 15,582 33.15 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
@@
HAT-F-Basis-First (Ours, ablation) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
HAT-F-Concat (Ours, ablation) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
@@ 5 Experiments
In this section, we provide the experiment setup and performance of Super-Resolution and Image Compression tasks. Please refer to Supplementary Materials for more ablation studies, training, implementation details, and analysis.
@@ 5.1 Image Super-Resolution
Experimental Setup. Following the same-task pretraining strategy of [15, 21], we pretrain all Super-Resolution (SR) models on ImageNet[23]. Our Factorized Features plug a shared Basis Transformer into four pretrained SR backbones—SwinIR [55], HAT [16], DAT [18], and ATD [110]. Backbones keep their original weights, while Basis Transformers start randomly. Pretraining runs 300k iterations on ImageNet with AdamW (lr 2e-4, batch 32, $\beta$ = 0.9/0.99); finetuning runs 200k iterations on DF2K (DIV2K [1]+Flickr2K [59]) with lr 1e-5. Inputs are $256\times 256$ crops, bicubically down-sampled to $64\times 64$ for the backbone. We set the number of coefficient–basis pairs to $N=6$ and use frequency scalars $\alpha_{j}\in\{1,4,16,64\}$ to capture both low- and high-frequency details.
Quantitative Results. table˜1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [6], Set14 [107], BSD100 [75], Urban100 [38], and Manga109 [76]. For quantitative metrics, PSNR and SSIM are reported. Our model delivers a 204.4% average PSNR gain, defined as (cb)/(ab)(c-b)/(a-b) with bb=SwinIR, aa=HAT-L, and cc=HAT-L-F. We embed three SoTA SR backbones: ATD [109], DAT [18], HAT [16], yielding ATD-F, DAT-F, and HAT-F. Each beats its parameter-matched baseline (ATD-L, DAT-L, HAT-L). As only HAT provides a large variant, we upscale ATD and DAT to the same size and train all models with identical pre-training and fine-tuning for fairness.
@@
The Order of Coefficient and Basis. Furthermore, in traditional Fourier Series and other image processing methods [95], the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in  fig.˜5a. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
@@
The Importance of Factorized Features Reconstruction. Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in eq.˜7. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized Features Reconstruction, highlighting its effectiveness.
@@
Visual Comparison. We provide the visual comparison in fig.˜6. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
@@
Method BD-Rate (%) $\downarrow$ Total Encoding Time (s) $\downarrow$ Total Decoding Time (s) $\downarrow$ Params(M)
@@
Cheng (CVPR’20)[20] 5.44 1.98 4.69 29.6
Xie (MM’21)[101] -0.78 2.93 6.00 50.0
STF (CVPR’22)[116] -4.31 0.14 0.13 99.9
ELIC (CVPR’22)[36] -7.24 0.07 0.09 36.9
TCM (CVPR’23)[62] -11.74 0.16 0.15 76.7
SegPIC (ECCV’24) [67] -8.26 0.14 0.13 83.5
LALIC (CVPR’25) [27] -15.26 0.27 0.15 63.2
@@
TCM-HAT-F-multi M=1 (Ours) 27.96 0.2320 0.1742 131.35
TCM-HAT-F-multi M=2 (Ours) 2.70 0.1998 0.1270 131.35
TCM-HAT-F-multi M=4 (Ours) -10.11 0.1846 0.1039 131.35
TCM-HAT-F-multi M=8 (Ours) -16.61 0.1756 0.0922 131.35
TCM-HAT-F-multi M=16 (Ours) -19.88 0.1715 0.0863 131.35
TCM-HAT-F-multi M=24 (Ours) -20.97 0.1702 0.0844 131.35